### v2

# Two-level Nonlinear Preconditioning Methods for Flood Models Posed on Perforated Domains

Abstract
This article focuses on the numerical solution of the Diffusive Wave equation posed in a domain containing a large number of polygonal perforations. These numerous perforations represent structures in urban areas, and this problem is used to model urban floods. This article relies on the work done in a previous article by the same authors, in which we introduced low-dimensional coarse approximation space for the linear Poisson equation based on a coarse polygonal partitioning of the domain. Similarly to other multi-scale numerical methods, this coarse space is spanned by locally discrete harmonic basis functions. We show that this coarse space also lends itself to the linearized Diffusive Wave problem that is obtained at each iteration of Newton’s method. Furthermore, we present nonlinear preconditioning techniques, including a Two-level RASPEN and Two-step method, which can significantly reduce the number of iterations compared to the traditional Newton’s method. For all proposed methods, we provide a discussion regarding the complexity of the algorithms. Numerical examples illustrating the performance of the proposed methods include a large-scale test case based on topographical data from the city of Nice.
journal: Computers and Mathematics with Applications
[…]
[uca]organization=Université Côte d’Azur, Laboratoire J.A. Dieudonné, CNRS UMR 7351,city=Nice, postcode=06108, country=France
[…]
[inria]organization=Team COFFEE, INRIA Centre at Université Côte d’Azur, city=Nice, postcode=06108, country=France
[…]
[tue]organization=Eindhoven University of Technology, Department of Mathematics and Computer Science, city=Eindhoven, postcode=5600 MB, country=The Netherlands
1 Introduction
In this article, we apply domain decomposition and multi-scale approaches to solve stationary and time-dependent partial differential equations. The main focus of the paper is the Diffusive Wave model [29, 2], given by
$\partial_{t}u-\text{div}\left(c_{f}h(u,z_{b}(\mathbf{x}))^{\alpha}||\nabla u||^{\gamma-1}\nabla u\right)=f,$
where $z_{b}(\mathbf{x})$ represents bathymetry/topography, $u$ is total water level, $h(u,z_{b}(\mathbf{x}))=\max(u-z_{b}(\mathbf{x},0)$ is the depth of the water with respect to bathymetry, $c_{f}$ denotes friction, $\gamma\leq 1$, and $1\leq\alpha\leq 2$. Generally, we take the source term $f$ as constant; this source term can be considered to represent rainfall. The Diffusive Wave Equation (DWE) is doubly nonlinear in the sense that there are nonlinearities both in the $(u-z_{b})$ term and in the gradient term. Additionally, the equation becomes degenerate, as $\kappa(u,\nabla u)\rightarrow 0$ as $h\rightarrow 0$ and $\kappa(u,\nabla u)\rightarrow\infty$ as $\nabla u\rightarrow 0$. We can think of the DWE as a simplification of the 2D Shallow Water Equations with zero-inertia assumptions ; for this reason, the DWE is often referred to as the Zero Inertia (ZI) equations [2]. This equation is commonly used to model overland flows.
With this model problem, we wish to model overland flows in urban areas, particularly the area of Nice, France. Anticipating these flood events can aid in the positioning of protective systems including dams, dikes, or rainwater drainage networks. One of the challenges of this numerical modeling of urban floods is that small structural features (buildings, walls, etc.) may significantly affect the water flow. In particular, the representation of the structural features in the numerical model impacts both the timing and extent of the flood [33, 34, 1]. The data set used in this article has been provided by Métropole Nice Côte d’Azur (MNCA) and allows for an infra-metric description of the urban geometries [3]. From the hydraulic perspective, these structural features can be assumed to be essentially impervious, and therefore represented as perforations (holes) in the model domain.
With this, let $D$ be an open, simply connected polygonal domain in $\mathbb{R}^{2}$. We denote by $\left(\Omega_{S,k}\right)_{k}$ a finite family of perforations in $D$ such that each $\Omega_{S,k}$ is an open connected polygonal subdomain of $D$. The perforations are mutually disjoint, that is $\overline{\Omega_{S,k}}\cap\overline{\Omega_{S,l}}=\emptyset$ for any $k\neq l$. We denote $\Omega_{S}=\bigcup_{k}\Omega_{S,k}$ and $\Omega=D\setminus\overline{\Omega_{S}}$, assuming that the family $\left(\Omega_{S,k}\right)_{k}$ is such that $\Omega$ is connected. The model problem of focus is given by the following: for a fixed time interval $(0,T]$,
$\left\{\begin{array}[]{rll}\displaystyle\partial_{t}u-\text{div}\left(\kappa(u,\nabla u)\nabla u\right)&=&0\qquad\mbox{in}\qquad\Omega\times(0,T],\\ \displaystyle u&=&g\qquad\mbox{on}\qquad(\partial\Omega\setminus\partial\Omega_{S})\times(0,T],\\ \displaystyle\kappa(u,\nabla u)\frac{\partial u}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad(\partial\Omega\cap\partial\Omega_{S})\times(0,T],\\ \end{array}\right.$ (1)
where $\mathbf{n}$ is the outward normal to the boundary, $\kappa(u,\nabla u)=h(u,z_{b}(\mathbf{x}))^{\alpha}||\nabla u||^{\gamma-1}$. The problem (1) must be combined with some initial condition on $u$ such that $u=u_{0}$ in $\Omega\times(t=0)$.
In the context of urban flood modeling, $\left(\Omega_{S,k}\right)_{k}$ can be thought of as a family of impervious structures such as buildings, fences, or other similar structures. However, problems posed on perforated domains also arise in multiple real-world scenarios, such as fluid/groundwater flow in porous media, atmospheric models, battery models, and drug delivery systems. After some discretization to be detailed later, this nonlinear equation can be written as the nonlinear system
$F(\mathbf{u})=0,$
where the numerical solution $\mathbf{u}$ approximates the exact solution $u$.
Depending on the geometrical complexity of the computational domain, the numerical solution of the PDE may become increasingly challenging. Our goal is to numerically solve the model problem on these perforated domains by using domain decomposition techniques. The latter use a divide and conquer approach to partition the domain into multiple smaller subdomains. Therefore, one has two “levels” of space discretization; the first level is based on a coarse polygonal partitioning of $\Omega$, while the second one is associated with the fine-scale triangulation and is designed to resolve the small-scale details of the model domain. The fine-scale triangulation is generated by Triangle [30].
Domain decomposition (DD) approaches to solve nonlinear PDEs can generally split into two categories: Newton-Krylov Schwarz (NKS) and Schwarz Newton-Krylov (SNK) methods. In NKS methods [6], the problem is first linearized, typically by Newton’s method or an inexact Newton method, which solve a linear system at each iteration. It is well-known that for large/complex linear systems, a direct linear solve can become quite expensive and time-consuming. In this case, one can then apply typical linear DD methods to this linear Newton system; i.e., use standard domain decomposition preconditioners coupled to a Krylov method on the linear system to be solved at each Newton iteration. Therefore, in the case of NKS methods applied to a nonlinear problem $F(\mathbf{u})=0$, the system remains unchanged, and we use DD methods to make the linear solve more efficient.
Stiff nonlinearities arise in many nonlinear problems, including problems such as heterogeneous problems with high-contrast coefficients. These stiff nonlinearities can cause a plateau in the norm of the Newton residual before/if the quadratic region of convergence is reached. In these cases in a global Newton’s method, one may be forced to take very small time steps, or damp the Newton iteration to an extremely small damping factor. To combat this, more recently, SNK methods were introduced; these methods do not begin with a linearization and are commonly referred to as nonlinear preconditioning methods. In SNK methods, the nonlinear system itself is changed, unlike in NKS methods. The nonlinear preconditioners involve the solution of local nonlinear problems and can therefore localize stiff nonlinearities.
Specifically, given a nonlinear system $F(\mathbf{u})=0$, these SNK preconditioning methods solve a new nonlinear system $\mathcal{F}(\mathbf{u})$ via Newton’s or inexact Newton’s method, where if $\mathcal{F}(\mathbf{u})=0$, then $F(\mathbf{u})=0$ (i.e., the new nonlinear system shares solutions with the original nonlinear system). This system $\mathcal{F}(\mathbf{u})=0$ is referred to as the nonlinearly preconditioned system. The latter is generated by applying Newton’s or inexact Newton’s method to a fixed point equation. The goal is for the nonlinearly preconditioned system to be solved faster than the original one; that is, to more quickly advance to the region of quadratic convergence for Newton’s method and reduce the number of exact or inexact Newton iterations.
In the Additive Schwarz Preconditioned Inexact Newton (ASPIN) [7], the authors introduce a method, where the nonlinearly preconditioned system $\mathcal{F}(u)=0$ consists of more balanced/uniform nonlinearities. ASPIN involves the solution of local nonlinear problems, using similar restriction and extension operators that are used in the linear case. Then, a global linear problem is solved via inexact Newton. In theory, for difficult problems with intense nonlinearities, this technique can accelerate convergence when compared to a typical Newton or inexact Newton method. Additionally, it can aid in the occurrence of stagnation in Newton’s method and desensitize Newton to the initial guess. We remark as well the existence of the MSPIN method [26], where generally, the partitioning of degrees of freedom is based on field type (field-splitting) instead of the typical subdomain partitioning. A main difference between the methods is that MSPIN can be thought of as analogous to the Multiplicative Schwarz (MS) method for linear systems, a generally non-parallel method, while ASPIN is analogous to the Additive Schwarz method (ASM). With this, one can also think of ASPIN and MSPIN as varying from each other similarly to Jacobi and Gauss-Seidel methods. In MSPIN and ASPIN, inexact Newton solves are computed for the nonlinearly preconditioned system.
The authors of [15] extended the ASPIN method to form Restricted Additive Schwarz Preconditioned Exact Newton (RASPEN), which can be thought of as analogous to the linear RAS method. Similarly to the linear case, RASPEN does not sum contributions in the overlap, using partition of unity operators to form a fixed point iteration to which Newton’s method can be applied. The authors of [15] showed that by design, the nonlinearly preconditioned system $\mathcal{F}_{RASPEN}(\mathbf{u})=0$ can be solved via exact Newton, as an equation can be derived for the exact Jacobian matrix $\nabla\mathcal{F}_{RASPEN}(\mathbf{u})$.
The authors of [11] introduced SRASPEN, which is obtained similarly by applying Newton’s method to the fixed point equation resulting from the nonlinear SRAS (Substructured Restricted Additive Schwarz) method. Unlike the aforementioned methods, the SRASPEN method works on nonoverlapping subdomains; these nonoverlapping domain decomposition methods are often referred to as substructuring methods. Within the same spirit, the authors of [12] introduced DNPEN, a nonlinear preconditioning method which applies Newton’s method to the Dirichlet-Neumann (DN) method, which is a substructuring method based on matching fluxes along subdomain interfaces. Additionally, the authors of [27] propose an adaptive nonlinear preconditioning technique which “turns off” the nonlinear preconditioning for outer Newtons where it is not necessary.
The aforementioned techniques can be considered as being left preconditioning methods, as they are solving a new tangential system $\mathcal{F}(\mathbf{u})=0$ by exact or inexact Newton. We also remark that there are a family of right nonlinear preconditioning techniques; these methods change the unknowns of the original system. We can give as an example the following works [22, 28, 23, 24, 25], where the authors introduced nonlinear preconditioners analogous to the Finite Element Tearing and Interconnecting (FETI), Finite Element Tearing and Interconnecting Dual-Primal (FETI-DP), and Balancing Domain Decomposition by Constraints (BDDC) methods. Additionally, the authors of [8] introduced the Nonlinear Elimination Preconditioned Inexact Newton (NEPIN), where “problem” components of the residual are implicitly eliminated and a resulting modified Newton direction is computed.
As is the case for linear DD methods, two-level methods (methods which include a coarse correction/level), are often needed for scalability in nonlinear preconditioning methods. Particularly, the number of GMRES iterations for the global tangential solve will not be scalable without a coarse correction. The coarse component allows for global communication across all subdomains. In the original article introducing RASPEN [15], the authors proposed a Full Approximation Scheme (FAS) to apply the coarse level. For MSPIN, a multiplicative coarse correction counterpart was introduced. In [19], the author proposes a linear coarse level for ASPIN and in [17], the authors provide numerical evidence of scalability of two-level variants of ASPIN and RASPEN. In particular, they propose multiple approaches of adding the coarse space correction (both additively and multiplicatively) based on Galerkin projections, and a numerical comparison of the methods is reported for different types of coarse spaces.
Our main contributions are as follows. In this paper, we rely on the work done in [4], in which we introduced an MsFEM coarse space for the linear Poisson problem posed on the perforated domains. Here, for nonlinear problems on perforated domains, we focus on left nonlinear preconditioning based on overlapping subdomains, coupled to techniques that use the coarse space proposed in [4] to provide scalability. We wish to stress that this article tackles the dual challenge of multiscale phenomena and double nonlinearity in the model problem; simultaneously addressing these points increases the complexity and significance of our approach. We propose the linear coarse space from [4] as a component of a two-level Restricted Additive Schwarz preconditioner for a linearized Newton system, which arises from the linearization of a nonlinear equation. Additionally, we use this coarse space as a component for multiple nonlinear methods, including a Two-level RASPEN method and a “Two-step” method originally introduced in [9] under the name NKS-RAS. While the coarse space of [4] was originally created based on the linear Poisson equation, we find that it is a fitting coarse space for nonlinear problems as well, and we use this in the later-mentioned two-level nonlinear preconditioning methods. The proposed methods also include the Anderson acceleration of a two-level method [31]. We provide a detailed complexity analysis of the proposed methods, discussing the benefits and downsides of each method. For these nonlinear methods applied to a time-dependent problem, we implement an adaptive local time-stepping method, which allows us to avoid a global time step reduction. In addition, this article serves as a numerical application of nonlinear preconditioning methods applied to nonlinear problems, containing multiple numerical examples with the main example being the doubly nonlinear Diffusive Wave model posed on complex domains containing multiple perforations.
The remainder of the paper is structured as follows. Section 2 provides a detailed description of the finite-element/ finite-volume hybrid discretization used for the Diffusive Wave Equation. Section 3 describes the coarse space from [4], which will be used in the two-level methods described in the article. Section 4 summarizes domain decomposition methods for linear problems, presenting the well-known linear Additive Schwarz and Restricted Additive Schwarz methods. Section 5 describes the proposed methods in detail and discusses the cost/complexity of each proposed algorithm. Section 6 provides numerical results, with experiments based on a porous medium equation posed on both an L-shaped domain and a large urban domain, followed by the Diffusive Wave equation posed on a large urban domain. Section 7 concludes with a summary and description of future work.
2 Discretization of Diffusive Wave equation
We now proceed with a discretization of the Diffusive Wave model (1). Let $T_{f}>0$ be the final flow simulation time and let $0=t_{0}<t_{1}<\ldots<t_{N_{T}}=T_{f}$ be a family of real numbers such that $\Delta t_{n}=t_{n+1}-t_{n}$.
We consider the following semi-implicit time discretization of (1): for $u^{n+1}=u|_{t_{n+1}}$,
$\left\{\begin{array}[]{rll}\displaystyle\frac{u^{n+1}-u^{n}}{\Delta t_{n}}-\text{div}\left(c_{f}\kappa(u^{n+1},\nabla u^{n})\nabla u^{n+1}\right)&=&0\qquad\mbox{in}\qquad\Omega,\\ \displaystyle u^{n+1}&=&g\qquad\mbox{on}\qquad\partial\Omega\setminus\partial\Omega_{S},\\ \displaystyle c_{f}\kappa(u^{n+1},\nabla u^{n})\frac{\partial u^{n+1}}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad\partial\Omega\cap\partial\Omega_{S},\end{array}\right.$ (2)
where $u^{0}$ is a provided initial condition.
Let $\mathcal{T}$ denote the triangulation of $\Omega$ which is conforming with respect to the partitioning $\left(\Omega_{j}\right)_{j=1}^{N}$. We introduce the finite element space defined by
$V_{h}=\{v\,\,|\,\,v\in C^{0}(\overline{\Omega}),\quad\text{s.t.}\quad v|_{t}\in\mathbb{P}_{1},\quad\text{for all}\quad t\in\mathcal{T}\}.$
Recalling the definition of the space $H^{1}_{\partial\Omega\setminus\partial\Omega_{S}}(\Omega)$ defined in (7), let $V_{h,0}=V_{h}\cap H^{1}_{\partial\Omega\setminus\partial\Omega_{S}}(\Omega)$ and let $g_{h}\in V_{h}$ be some approximation of the boundary data $g$. The finite element discretization of (2) reads as: Find $u_{h}^{n+1}\in V_{h}$ satisfying $u_{h}^{n+1}|_{\partial\Omega\setminus\Omega_{S}}=g_{h}$ such that
$\frac{1}{\Delta t_{n}}\int_{\Omega}(u_{h}^{n+1}-u_{h}^{n})\,v_{h}\,\mathrm{d}\mathbf{x}+\int_{\Omega}c_{f}\kappa(u_{h}^{n+1},\nabla u_{h}^{n})\nabla u_{h}^{n+1}\cdot\nabla v_{h}\,\mathrm{d}\mathbf{x}=0\qquad\mbox{for all}\quad v_{h}\in V_{h,0}.$ (3)
Let $(\mathbf{x}_{i})_{i=1}^{N_{\Omega}}$ denote the set of triangulation points; the set of point indices $\{1,2,...,N_{\Omega}\}$ is denoted by $\mathcal{N}$. We denote the nodal “hat” basis functions $(\eta^{\ell})_{\ell=1}^{N_{\Omega}}$. Furthermore, we define $u_{i}^{n+1}$ and $u_{\ell}^{n+1}$ as the nodal values of the discrete solution at nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{\ell}$ at $t_{n+1}$, respectively. Additionally, we set ${\cal N}_{T}=\{i\in{\cal N}\,\,|\,\,\mathbf{x}_{i}\in\overline{T}\}$ for any triangle $T\in{\cal T}$ ($T$ assumed to be open) and denote $\mathcal{T}_{i}$ as the subset of triangles connected to the node $i$ such that $\mathcal{T}_{i}=\{T\in\mathcal{T}\,\,|\,\,\mathbf{x}_{i}\in\overline{T}\}$. Furthermore, we denote by ${\cal N}_{i}$ the set of notes adjacent to node $i$ such that ${\cal N}_{i}=\bigcup_{T\in\mathcal{T}_{i}}{\cal N}_{T}.$ Finally, let ${\cal N}_{D}$ denote the set of Dirichlet nodes such that ${\cal N}_{D}=\{i\in{\cal N}\,\,|\,\,\mathbf{x}_{i}\in\partial\Omega\setminus\Omega_{S}\}$.
In order to enhance the stability of the numerical scheme, we modify the finite element scheme by means of discretization techniques from finite volume methods. In particular, we introduce mass lumping in the accumulation and upwinding in the diffusion term. This discretization technique is generally referred to as a Finite Volume- Finite Element (FV-FE) discretization [13], or it is often referred to as a Control Volume Finite Element method [10, 14].
Diffusion term: Since $\sum_{\ell\in{\cal N}_{T}}\nabla\eta_{\ell}=0$ on each triangle $T$, we observe for the diffusive term that for $i\in{\cal N}$,
[…]
In order to deal with the degeneracy of $\kappa(u,\boldsymbol{\xi})=h(u,z_{b})^{\alpha}\|\boldsymbol{\xi}\|^{\gamma-1}$ for $h(u,z_{b})=0$, we introduce upwinding in our discretization. Denoting
$\tau_{i\ell,T}^{n}=-c_{f}\Bigl{|}\nabla u_{h}^{n}|_{T}\Bigr{|}^{1-\gamma}\int_{T}\nabla\eta_{\ell}\cdot\nabla\eta_{i}\,{\rm d}\mathbf{x}\qquad\mbox{and}\qquad\tau_{i\ell}^{n}=\sum_{T\in{\cal T}_{i}\cap{\cal T}_{\ell}}\tau_{i\ell,T}^{n},$
we introduce upstream water depth defined by
$h_{i\ell}^{n+1}=\left\{\begin{array}[]{llll}h(u^{n+1}_{i},z_{b}({\bf x}_{i}))&\mbox{if }\tau_{i\ell}^{n}(u^{n+1}_{i}-u_{\ell}^{n+1})\geq 0,\\ h(u_{\ell}^{n+1},z_{b}({\bf x}_{\ell}))&\mbox{otherwise}.\end{array}\right.$
Then, we approximate
$\int_{\Omega}c_{f}\kappa(u_{h}^{n+1},\nabla u_{h}^{n})\nabla u_{h}^{n+1}\cdot\nabla\eta_{i}\,{\rm d}\mathbf{x}\approx\sum_{\ell\in{\cal N}_{i}}\tau_{i\ell}^{n}\left(h_{i\ell}^{n+1}\right)^{\alpha}(u_{i}^{n+1}-u_{\ell}^{n+1})\qquad\mbox{for all}\quad i\in{\cal N}\setminus\mathcal{N}_{D}.$
Accumulation term: Let $M_{i\ell}$ denote the $(i,\ell)$th entry of a classical finite element matrix, that is ${M_{i\ell}=\int_{\Omega}\eta_{i}\eta_{\ell}\,\,{\rm d}\mathbf{x}}.$ Then, denoting $m_{i}=\sum_{\ell=1}^{N_{\Omega}}M_{i\ell},$ we approximate
$\frac{1}{\Delta t_{n}}\int_{\Omega}(u_{h}^{n+1}-u_{h}^{n})\cdot\,\eta_{i}\,\mathrm{d}\mathbf{x}$ $\approx\frac{m_{i}}{\Delta t_{n}}\left(u_{i}^{n+1}-u_{i}^{n}\right)\qquad\mbox{for all}\quad i\in{\cal N}\setminus\mathcal{N}_{D}.$
The resulting combined FV-FE scheme writes as
$\begin{cases}\displaystyle\frac{m_{i}}{\Delta t_{n}}\left(u_{i}^{n+1}-u_{i}^{n}\right)+\sum_{\ell\in{\cal N}_{i}}\tau_{i\ell}^{n}\left(h_{i\ell}^{n+1}\right)^{\alpha}(u_{i}^{n+1}-u_{\ell}^{n+1})=0\quad&\text{for}\quad i\in{\cal N}\setminus\mathcal{N}_{D},\\ \displaystyle u_{i}^{n+1}=g_{h}(\mathbf{x}_{i}),\quad&\text{for}\quad i\in\mathcal{N}_{D}.\end{cases}$ (4)
In matrix-vector notations for a given vector $\mathbf{u}=(u_{i})_{i\in{\cal N}\setminus{\cal N}_{D}}$, the scheme can be summarized as follows: for time step $n=0,1,\ldots,$ solve
$F(\mathbf{u}^{n+1})=0,$
where the $i$th component of $F(\mathbf{u}^{n+1})$ is given by
$F(\mathbf{u}^{n+1})_{i}=\frac{m_{i}}{\Delta t_{n}}\left(u_{i}^{n+1}-u_{i}^{n}\right)+\sum_{\ell\in\mathcal{N}_{i}}\tau_{i\ell}^{n}\left(h_{i\ell}^{n+1}\right)^{\alpha}(u_{i}^{n+1}-u_{\ell}^{n+1})\quad\text{for }i\in\mathcal{N}\setminus\mathcal{N}_{D},$ (5)
and $u_{i}^{n+1}=g_{h}(\mathbf{x}_{i})$ for $i\in\mathcal{N}_{D}$.
3 Multiscale Coarse space
We now briefly describe the coarse space that will be used for all introduced two-level algorithms; we refer the reader to [4] for a detailed description. The coarse space is of the multiscale type, composed of piecewise harmonic basis functions with respect to a polygonal partitioning of the domain. The latter are computed via the numerical solutions of linear Laplace problems.
Consider a finite nonoverlapping polygonal partitioning of $D$ denoted by $\left(D_{j}\right)_{j=1,\ldots,N}$ and an induced nonoverlapping partitioning of $\Omega$ denoted by $\left(\Omega_{j}\right)_{j=1,\ldots,N}$ such that $\Omega_{j}=D_{j}\cap\Omega$. We will refer to $\left(\Omega_{j}\right)_{j=1,\ldots,N}$ as the coarse mesh over $\Omega$. Additionally, we denote by $\Gamma$ its skeleton, that is $\Gamma=\bigcup_{j=1,\ldots,N}\partial\Omega_{j}\setminus\partial\Omega_{S}$. Let $\left(e_{k}\right)_{k=1,\ldots,N_{e}}$ denote a nonoverlapping partitioning of the $\Gamma$ such that each “coarse edge” $e_{k}$ is an open planar segment, and we denote $H=\max_{k=1,\ldots,N_{e}}|e_{k}|$. The set of coarse grid nodes is given by ${\cal V}=\bigcup_{k=1,\ldots,N_{e}}\partial e_{k}$. Figure 1 illustrates the location of the coarse grid nodes that typically result from clipping $(D_{j})_{j}$ with $\Omega_{S}$.
Figure 1: Coarse cell $\Omega_{j}$, nonoverlapping skeleton $\Gamma$ (blue lines), and coarse grid nodes $\mathbf{x}_{s}\in{\cal V}$ (red dots). Coarse grid nodes are located at $\overline{\Gamma}\cap\partial\Omega_{S}.$
Let $H^{1}_{\Delta}(\Omega)$ be a subspace of $H^{1}(\Omega)$ composed of piece-wise harmonic functions, weakly satisfying the homogeneous Neumann boundary conditions on $\partial\Omega\cap\partial\Omega_{S}$, that is
$H_{\Delta}^{1}(\Omega)=\{\,u\in H^{1}(\Omega)\,\,|\,\,(u|_{\Omega_{j}},v)_{H^{1}(\Omega_{j})}=0\quad\text{for all}\quad v\in H^{1}_{\partial\Omega\setminus\partial\Omega_{S}}(\Omega_{j})\},$ (6)
where
$H^{1}_{\partial\Omega\setminus\partial\Omega_{S}}(\Omega)=\{u\in H^{1}(\Omega)\,|\,u|_{\partial\Omega\setminus\partial\Omega_{S}}=0\}.$ (7)
In other words, $u\in H_{\Delta}^{1}(\Omega)$ if and only if $u\in H^{1}(\Omega)$ and for all subdomains $\Omega_{j}$, the equations
$\left\{\begin{array}[]{rll}-\Delta u|_{\Omega_{j}}&=&0\qquad\mbox{in}\qquad\Omega_{j},\\ \displaystyle\frac{\partial u}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad\partial\Omega_{j}\cap\partial\Omega_{S},\\ \end{array}\right.$ (8)
are satisfied in a weak sense.
With this, we define
$V_{H}^{\Gamma}=\{\,v\in C^{0}(\overline{\Gamma})\,\,|\,\,v|_{e_{k}}\in\mathbb{P}_{1}(e_{k})\quad\text{for all}\quad k=1,\ldots,N_{e}\},$
where $\mathbb{P}_{1}(e_{k})$ denotes the set of piecewise linear polynomials over an edge $e$. We also define
$V_{H}=\{\,v\in H_{\Delta}^{1}(\Omega)\,\,|\,\,v|_{\Gamma}\in V_{H}^{\Gamma}\}.$
Let $(g_{s})_{s=1,\ldots,N_{{\cal V}}}$ be the nodal basis of $V_{H}^{\Gamma}$, where $N_{\cal V}$ denotes the total number of coarse grid nodes. In practice, $(g_{s})_{s=1,\ldots,N_{{\cal V}}}$ is composed of nodal piecewise linear “hat” functions. The function $\phi_{s}\in V_{H}$ associated to $g_{s}$ is computed by weakly imposing on each $\Omega_{j}$
$\left\{\begin{array}[]{rllll}\Delta\phi_{\bf s}&=&0&\mbox{in}&\Omega_{j},\\ \displaystyle\frac{\partial\phi_{s}}{\partial\bf{n}}&=&0&\mbox{on}&\partial\Omega_{j}\cap\partial\Omega_{S},\\ \phi_{s}&=&g_{s}&\mbox{on}&\partial\Omega_{j}\setminus\partial\Omega_{S}.\\ \end{array}\right.$ (9)
We refer to the space $V_{H}$ as the Trefftz space. The basis of the discrete version of the Trefftz space is obtained through the finite element approximation of (9) which results in the system of the form
$\tilde{A}_{j}\boldsymbol{\phi}_{s}^{j}=b_{s}^{j},$
where $\tilde{A}_{j}$ is the local stiffness matrix and $b_{s}^{j}$ accounts for the Dirichlet boundary data in (9).
Let $\overline{R}_{j}$ denote the boolean restriction matrices which restrict from the degrees of freedom in the discretization of $\Omega$ to the degrees of freedom in the nonoverlapping subdomain $\overline{\Omega_{j}}$, and let $\overline{D}_{j}$ denote partition of unity matrices corresponding to the discretization of $\overline{\Omega_{j}}$ such that $\sum_{j=1}^{N}R_{j}^{T}D_{j}R_{j}=I$. Furthermore, let $\boldsymbol{\phi}_{s}$ be a global vector which, when restricted to a subdomain $\Omega_{j}$, returns $\boldsymbol{\phi}_{s}^{j}$. That is,
$\boldsymbol{\phi}_{s}=\sum_{j\in\mathcal{N}_{s}}\overline{R}_{j}^{T}\overline{D}_{j}\boldsymbol{\phi}_{s}^{j},$
for $s=1,\ldots,N_{\cal V}$, where $\mathcal{N}_{s}=\{\,j\,\,|\,\,\mathbf{x}_{s}\,\,\text{is contained in}\,\,\Omega_{j}\,\}.$ The global vectors $\boldsymbol{\phi}_{s}$ are the basis vectors associated to the coarse grid nodes $\mathbf{x}_{s}\in{\cal V}$. The discrete Trefftz space is then defined as the span of the basis vectors $\boldsymbol{\phi}_{s},s=1,\ldots,N_{{\cal V}}$. With this, the coarse transition matrix $R_{H}$ is such that the $k$th row of $R_{H}$ is given by $\boldsymbol{\phi}_{k}^{T}$ for $k=1,\ldots,N_{\cal V}$.
4 Linear Multi-domain Solution Methods
In this section, we summarize linear DD methods, particularly overlapping Schwarz methods, that will be extended in the next section to nonlinear PDEs. Consider some discretized linear system of the form
$A\mathbf{u}=\mathbf{f}.$
Let $\left(\Omega_{j}^{\prime}\right)_{j=1,\ldots,N}$ denote the overlapping partitioning of $\Omega$ such that $\Omega_{j}\subset\Omega_{j}^{\prime}$. In practice, each $\Omega_{j}^{\prime}$ is constructed by propagating $\Omega_{j}$ by a few layers of triangles. With this, let $R_{j}$ denote the boolean restriction matrices which restrict from the degrees of freedom in the discretization of $\Omega$ to the degrees of freedom in the overlapping subdomain $\Omega_{j}^{\prime}$, and let $D_{j}$ denote partition of unity matrices corresponding to the discretization of $\Omega_{j}^{\prime}$ such that $\sum_{j=1}^{N}R_{j}^{T}D_{j}R_{j}=I$.
With this, the iterative RAS method computes the approximation to the solution $\mathbf{u}^{k+1}$ from $\mathbf{u}^{k}$ by
$\begin{array}[]{llll}\mathbf{u}^{k+1}&=\mathbf{u}^{k}+M_{RAS,1}^{-1}(\mathbf{f}-A\mathbf{u}^{k}),\\ \end{array}$ (10)
where
$M_{RAS,1}^{-1}=\sum_{j=1}^{N}R_{j}^{T}D_{j}[R_{j}AR_{j}^{T}]^{-1}R_{j}.$
Additionally, we can accelerate this fixed-point method (10) by solving the preconditioned system
$M_{RAS,1}^{-1}A\mathbf{u}=M_{RAS,1}^{-1}{\bf f},$ (11)
with a Krylov type method such as GMRES.
Generally, the number of iterations in (10) and (11) grows with the number of subdomains in one direction. For example, this increase is linear for the stationary iteration (10). To combat this, a coarse level is needed to aid in global communication between all subdomains. Consider the coarse matrix $R_{H}$ defined in Section 3; with this, the two-level preconditioned Krylov method is given by
$M_{RAS,2}^{-1}A\mathbf{u}=M_{RAS,2}^{-1}{\bf f},$ (12)
where
$M_{RAS,2}^{-1}=R_{H}^{T}(R_{H}AR_{H}^{T})^{-1}R_{H}+\sum_{j=1}^{N}R_{j}^{T}D_{j}[R_{j}AR_{j}^{T}]^{-1}R_{j}.$ (13)
5 Nonlinear Multi-domain Solution Methods
We present results for various methods to solve the nonlinear system $F(\mathbf{u})=0$, where $F(\mathbf{u})$ arises from some discretization of a nonlinear PDE. For the Diffusive Wave model, the residual function $F(\mathbf{u})=0$ is given in (5).
Preconditioning the linearized Newton system
To solve nonlinear equations, Newton’s method is commonly employed. For exact Newton’s method, one needs the residual $F(\mathbf{u})$ as well as the Jacobian $\nabla F(\mathbf{u})$. The exact Newton’s method is described in Algorithm 1.
Algorithm 1 Newton’s method to solve $F(\mathbf{u})=0$
$\mathbf{u}^{0}$, residual function $F$, Jacobian $\nabla F$
for outer iteration $k=0,\ldots,$ to convergence do
     Compute Newton update $\boldsymbol{\delta}=[\nabla F(\mathbf{u})]^{-1}F(\mathbf{u})$
     Compute $\mathbf{u}^{k+1}=\mathbf{u}^{k}-\boldsymbol{\delta}$;
end for
Often when implementing Newton’s method, a backtracking line search is implemented to control the step size of Newton’s method. This often aids in convergence for certain problems, as the line search prevents “overshooting” in the Newton step direction. Newton’s method with backtracking line search is given in Procedure 1.
Procedure 1 Newton’s Method with backtracking line search to solve $F(\mathbf{u})=0$
for outer iteration $k=0,\ldots,$ to convergence do
     $d=1$
     Compute Newton update $\boldsymbol{\delta}=[\nabla F(\mathbf{u})]^{-1}F(\mathbf{u})$
     while $||F(\mathbf{u}-d\boldsymbol{\delta})||>(1-\frac{d}{4})||F(\mathbf{u})||$ do
         $d=\frac{d}{2}$
     end while
     Compute $\mathbf{u}^{k+1}=\mathbf{u}^{n}-d\,\boldsymbol{\delta}$;
end for
When a linear system is very large, a direct solve is often too expensive and not recommended. Therefore, it is common to solve the linear system in Algorithm 1 via a preconditioned Krylov solver. We proposed in [4] and recalled in Section 3 a coarse space for the two-level RAS preconditioner applied to the Poisson equation on perforated domains. It turns out that the same coarse space can be successfully employed within the two-level RAS preconditioner (13) for the linear system in Algorithm 1. The numerical evidences regarding the efficiency of such a two-level solver are presented in Section 6.
Let $J_{k}=\nabla F(\mathbf{u}^{k})$ denote the Jacobian matrix and $\mathbf{F}_{k}=F(\mathbf{u}^{k})$ the residual at a given iterate. The Newton’s method update denoted by $\boldsymbol{\delta}_{k}$ satisfies the system
$J_{k}\boldsymbol{\delta}_{k}=\mathbf{F}_{k}.$
The preconditioned system to solve for $\boldsymbol{\delta}_{k}$ is as follows:
$M_{RAS,2}^{-1}J_{k}\boldsymbol{\delta}_{k}=M_{RAS,2}^{-1}\mathbf{F}_{k},$
where $M_{RAS,2}$ is defined as in (13) with system matrix $A$ replaced by $J_{k}$,
$M_{RAS,2}^{-1}=\sum_{j=1}^{N}R_{j}^{T}D_{j}(R_{j}J_{k}R_{j}^{T})^{-1}R_{j}+R_{H}^{T}(R_{H}J_{k}R_{H}^{T})^{-1}R_{H}.$ (14)
Nonlinear Restrictive Additive Schwarz Iteration
Recall for linear systems $A\mathbf{u}={\bf f}$, the Restricted Additive Schwarz (RAS) method can be used as an iterative method or as a preconditioner for Krylov methods. Likewise, for the nonlinear problem, we introduce a nonlinear RAS (NRAS) fixed point iteration. Multiple methods which are to be introduced will require this nonlinear RAS iteration.
Definition 5.1 (An NRAS iteration).
Let $R_{j}$ be the boolean restriction matrices corresponding to the degrees of freedom of $\Omega_{j}^{\prime}$. Let the “subdomain update” function $G_{j}$, such that, for all $\mathbf{u}$, $G_{j}(\mathbf{u})$ is the solution of the local nonlinear system
$R_{j}F(R_{j}^{T}G_{j}(\mathbf{u})+(I-R_{j}^{T}R_{j})\mathbf{u})=0.$ (15)
An NRAS iteration is computed by first solving each (15) then glue local contributions:
$\text{NRAS}(\mathbf{u})=\sum_{j}R_{j}^{T}D_{j}G_{j}(\mathbf{u}).$ (16)
We can use this method as a fixed point iteration, computing iteratively the solution $\bf{u}^{k}$ starting from an initial guess $\bf{u}^{0}$:
$\mathbf{u}^{k+1}=\text{NRAS}(\mathbf{u}^{k}).$ (17)
Note that the subdomain updates $G_{j}(\mathbf{u})$ correspond to solving the local problem on $\Omega_{j}^{\prime}$ with Dirichlet boundary conditions. These problems (15) are solved via Newton’s method; as the subproblems are small, each Newton linear system is solved using a direct solver.
Remark 5.1 (NRAS is RAS when applied to a linear problem).
We comment here on the connection between the nonlinear RAS fixed point iteration (17) and the linear RAS fixed point iteration (10). Consider the local solutions computed from (15). For linear problems such that $F(\mathbf{u})=A\mathbf{u}-\mathbf{f}$, we have
$R_{j}F(R_{j}^{T}$ $G_{j}(\mathbf{u})+(I-R_{j}^{T}R_{j})\mathbf{u})=0$ (18)
$R_{j}(A(R_{j}^{T}$ $G_{j}(\mathbf{u})+(I-R_{j}^{T}R_{j})\mathbf{u})-\mathbf{f})=0$
$R_{j}AR_{j}^{T}$ $G_{j}(\mathbf{u})=R_{j}AR_{j}^{T}R_{j}\mathbf{u}+R_{j}(\mathbf{f}-A\mathbf{u})$
$G_{j}(\mathbf{u})=R_{j}\mathbf{u}+(R_{j}AR_{j}^{T})^{-1}R_{j}(\mathbf{f}-A\mathbf{u})$
such that the fixed point iteration (17) would be equivalent to
$\mathbf{u}^{k+1}$ $=\sum_{j=1}^{N}R_{j}^{T}D_{j}\left(R_{j}\mathbf{u}^{k}+(R_{j}AR_{j}^{T})^{-1}R_{j}(\mathbf{f}-A\mathbf{u}^{k})\right)$ (19)
$=\mathbf{u}^{k}+\sum_{j=1}^{N}R_{j}^{T}D_{j}(R_{j}AR_{j}^{T})^{-1}R_{j}(\mathbf{f}-A\mathbf{u}^{k})$
$=\mathbf{u}^{k}+M_{RAS,1}^{-1}(\mathbf{f}-A\mathbf{u}^{k}),$
where we have used the partition of unity property $I=\sum_{j=1}^{N}R_{j}^{T}D_{j}R_{j}$. Therefore, the NRAS fixed point iteration (17) reduces to the linear RAS fixed point iteration (10).
One-level RASPEN
Recall that in the linear case, we can accelerate the fixed point RAS iteration by Krylov methods. Analogously, in the nonlinear case, we can accelerate the NRAS fixed point iteration (17) by applying Newton’s method to the fixed point equation
$\mathcal{F}(\mathbf{u}):=\mathbf{u}-\text{NRAS}(\mathbf{u})=0$
resulting in the One-level RASPEN method [15] detailed in Algorithm 2. It was shown in [15] that the Jacobian of $\mathcal{F}(\mathbf{u})$ can be computed directly.
Algorithm 2 One-level RASPEN to solve $F(\mathbf{u})=0$
$\mathbf{u}^{0}$, residual function $F$, Jacobian $\nabla F$
for outer iteration $k=0,\ldots,$ to convergence do
     Compute NRAS update $\widehat{\mathbf{u}}^{k}=$ NRAS($\mathbf{u}^{k}$);
     Set $\mathcal{F}(\mathbf{u}^{k})=\mathbf{u}^{k}-\widehat{\mathbf{u}}^{k}$
     Solve $\mathbf{u}^{k+1}=\mathbf{u}^{k}-[\nabla\mathcal{F}(\mathbf{u}^{k})]^{-1}\mathcal{F}(\mathbf{u}^{k})$.
end for
In other words, solve $\mathcal{F}(\mathbf{u})=0$ via Newton’s method.
Remark 5.2 (Exact computation of the Jacobian).
Unlike in the case of ASPIN, for RASPEN the Jacobian is computed exactly. We describe here the computation of the Jacobian $\nabla\mathcal{F}(\mathbf{u}^{k})$ for Algorithm 2. From (16), we have
$\mathcal{F}(\mathbf{u}^{k})$ $=\mathbf{u}^{k}-\text{NRAS}({\mathbf{u}}^{k})=\mathbf{u}^{k}-\sum_{j}R_{j}^{T}D_{j}G_{j}(\mathbf{u}^{k}).$
Then the Jacobian is given by
$\nabla\mathcal{F}(\mathbf{u}^{k})$ $=\nabla\left(\mathbf{u}^{k}-\sum_{j}R_{j}^{T}D_{j}G_{j}(\mathbf{u}^{k})\right)=I-\sum_{j}R_{j}^{T}D_{j}\nabla G_{j}(\mathbf{u}^{k}).$ (20)
To evaluate this, we must compute $\nabla G_{j}(\mathbf{u}^{k})$ by the gradient of (15):
$\nabla G_{j}(\mathbf{u}^{k})=R_{j}-[R_{j}\nabla F(\mathbf{v}_{j}^{k})R_{j}^{T}]^{-1}R_{j}\nabla F(\mathbf{v}_{j}^{k})$ (21)
where we have set $\mathbf{v}_{j}^{k}=R_{j}^{T}G_{j}(\mathbf{u}^{k})+(I-R_{j}^{T}R_{j})\mathbf{u}^{k}$ for convenience. Thus, we have
$\nabla\mathcal{F}(\mathbf{u}^{k})$ $=I-\sum_{j}R_{j}^{T}D_{j}\nabla G_{j}(\mathbf{u}^{k}),$
$=I-\sum_{j}R_{j}^{T}D_{j}(R_{j}-[R_{j}\nabla F(\mathbf{v}_{j}^{k})R_{j}^{T}]^{-1}R_{j}\nabla F(\mathbf{v}_{j}^{k})),$
$=\sum_{j}R_{j}^{T}D_{j}[R_{j}\nabla F(\mathbf{v}_{j}^{k})R_{j}^{T}]^{-1}R_{j}\nabla F(\mathbf{v}_{j}^{k}),$ (22)
where we have used the partition of unity property $I=\sum_{j}R_{j}^{T}D_{j}R_{j}$. We note that the matrix $[R_{j}\nabla F(\mathbf{v}_{j}^{k})R_{j}^{T}]^{-1}$ can be dense and does not need to be directly assembled. Therefore, we set up $\nabla\mathcal{F}(\mathbf{u}^{k})$ as a linear operator, allowing us to compute the matrix-vector product $\nabla\mathcal{F}(\mathbf{u}^{k})x^{k}$ for any vector $x^{k}$ without explicitly computing the matrix $\nabla\mathcal{F}(\mathbf{u}^{k})$.
Two-level RASPEN
It is well-known that one-level domain decomposition methods are not scalable with respect to the number of subdomains. Specifically, the GMRES method used at each outer iteration of Algorithm 2 is not scalable without a coarse level. Therefore, we implement a Two-level RASPEN variant.
We choose a Two-level RASPEN correction provided in Algorithm 11 of [21], which allows the coarse correction to be done algebraically. The coarse correction is chosen to be computed after the local nonlinear solves. The authors of [17] provide numerical experiments on multiple different choices of two-level methods, including adding the coarse correction before, after, and before and after the nonlinear RAS iteration. The multiplicative Two-level RASPEN method we have chosen is given by Algorithm 3. This method solves the nonlinearly preconditioned system
$\mathcal{F}(\mathbf{u})=\mathbf{u}-\text{NRAS}(\mathbf{u})+R_{H}^{T}c_{H}(\text{NRAS}(\mathbf{u})),$ (23)
where the function $c_{H}(\mathbf{v})$ for a given $\mathbf{v}$ is defined as the solution to the “coarse” nonlinear equation
$R_{H}F(\mathbf{v}-R_{H}^{T}c_{H}(\mathbf{v}))=0.$ (24)
Algorithm 3 Two-level RASPEN to solve $F(\mathbf{u})=0$
$\mathbf{u}^{0}$, residual function $F$, Jacobian $\nabla F$
for outer iteration $k=0,\ldots,$ to convergence do
     Compute NRAS update $\widehat{\mathbf{u}}^{k}=\text{NRAS}(\mathbf{u}^{k})$;
     Solve coarse problem $R_{H}F(\widehat{\mathbf{u}}^{k}-R_{H}^{T}c_{H}(\widehat{\mathbf{u}}^{k}))=0$ for $c_{H}(\widehat{\mathbf{u}}^{k})$ via Newton’s method;
     Set $\mathcal{F}(\mathbf{u}^{k})=\mathbf{u}^{k}-\widehat{\mathbf{u}}^{k}+R_{H}^{T}c_{H}(\widehat{\mathbf{u}}^{k});$
     Solve $\mathbf{u}^{k+1}=\mathbf{u}^{k}-[\nabla\mathcal{F}(\mathbf{u}^{k})]^{-1}\mathcal{F}(\mathbf{u}^{k})$.
end for
In other words, solve $\mathcal{F}(\mathbf{u})=0$ via Newton’s method.
Two-step method
Next, we consider a Two-step method proposed in [9] under the name NKS-RAS; see also [5]. This method alternates between an NRAS iteration and a global Newton step at each outer iteration. This method leads to fast convergence (typically quadratic convergence is expected), but unlike Algorithms 2 and 3, can not be framed as a nonlinear preconditioning method. The Two-step method is given in Algorithm 4. The main feature of the Two-step method is the ease of its implementation. Compared to the standard Newton’s method, it simply requires an additional NRAS step; moreover, compared to the One and Two-level RASPEN methods, the Two-step method does not involve any dense linear systems, and therefore, in principle, may be implemented based on a direct linear solver. Here, the linearized system of Algorithm 4 is solved using the same preconditioned GMRES method as in Algorithm 1.
Algorithm 4 Two-step method to solve $F(\mathbf{u})=0$
$\mathbf{u}^{0}$, residual function $F$, Jacobian $\nabla F$
for outer iteration $k=0,\ldots,$ to convergence do
     Compute NRAS update $\widehat{\mathbf{u}}^{k}=$NRAS$(\mathbf{u}^{k})$;
     Compute Newton update $\mathbf{u}^{k+1}=\widehat{\mathbf{u}}^{k}-[\nabla F(\widehat{\mathbf{u}}^{k})]^{-1}F(\widehat{\mathbf{u}}^{k})$;
end for
Anderson Acceleration of Coarse Two-step Method
The Two-step method given in Algorithm 4 involved a global fine-scale linear solve at each outer iteration. We now propose to replace this step with a “coarse” linear solve, using the Trefftz coarse space described in Section 3. For this, we can employ a method similar to the Two-step method (4) but with a “coarse” Newton step at each iteration, in a method similar to our previous works on linear models. We refer to this as the Coarse Two-Step method, which allows us to use the coarse Trefftz space in the iteration and cheapen the cost of the global Newton update. This Coarse Two-step method is a fixed point iteration which involves a coarse Newton update, which we denote by
$F_{c}(\mathbf{u})=\mathbf{u}-R_{H}^{T}\boldsymbol{\delta}_{H},\quad\text{where}\quad\boldsymbol{\delta}_{H}=[R_{H}\nabla F(\mathbf{u})R_{H}^{T}]^{-1}R_{H}F(\mathbf{u}).$ (25)
With this, the fixed point coarse Two-step method is given as follows: for iteration $k=0,1,\ldots,$ to convergence, compute
$\widehat{\mathbf{u}}^{k}=\text{NRAS}(\mathbf{u}^{k}),$ (26)
$\mathbf{u}^{k+1}=F_{c}(\widehat{\mathbf{u}}^{k}).$
The convergence of the fixed point method (26) to the numerical solution is generally quite slow when compared to the original Two-step method given by Algorithm 4, as a full update is not done at each iteration. To combat this, we introduce Anderson Acceleration [31] as a way to accelerate the fixed point method (26), resulting in Algorithm 5. The general Anderson acceleration method is given in Procedure 2. Given some $\mathbf{u}^{k+1}=P(\mathbf{u}^{k})$, we solve $\mathcal{F}(\mathbf{u})=P(\mathbf{u})-\mathbf{u}=0$ via Anderson acceleration. While the parameter $m_{k}$ in Algorithm 2 can vary, we generally choose $m_{k}=10$ in our numerical experiments and do not explore this parameter further. For convenience, we will refer to the Anderson Acceleration of the Coarse Two-step method (Algorithm 5) as the “Anderson method” for the duration of the article. We note that Anderson’s method generally fails to accelerate quadratically convergent fixed point iterations; this is proven with an error estimate in [16].
Algorithm 5 Anderson Acceleration Applied to Coarse Two-step method to solve $F(\mathbf{u})=0$
$\mathbf{u}^{0}$, residual function $F$, Jacobian $\nabla F$
Denote $F_{c}(\mathbf{u}):=\mathbf{u}-R_{H}^{T}[R_{H}\nabla F(\mathbf{u})R_{H}^{T}]^{-1}R_{H}F(\mathbf{u})$;
Solve $V(\mathbf{u}):=F_{c}(\text{NRAS}(\mathbf{u}))=0$ by Anderson acceleration, (Procedure 2).
The solution to $V(\mathbf{u})=0$ is the solution to $F(\mathbf{u})=0$.
Procedure 2 Anderson Acceleration to solve some general function of the form $V(\mathbf{u})=P(\mathbf{u})-\mathbf{u}=0$
$\mathbf{u}^{0}$ and $m\geq 1$
Set $\mathbf{u}^{1}=P(\mathbf{u}^{0})$;
for outer iteration $k=1,2,\ldots,$ to convergence do
     Set $m_{k}=\min\{m,k\}$;
     Set $f_{k}=(\mathbf{V}_{k-m_{k}},...,\mathbf{V}_{k})$, where $\mathbf{V}_{i}=P(\mathbf{u}^{i})-\mathbf{u}^{i}$;
     Determine $\alpha^{k}=(\alpha_{0}^{k},\ldots,\alpha_{m_{k}}^{k})^{T}$ that solves
$\min_{\alpha^{k}=(\alpha_{0}^{k},\ldots,\alpha_{m_{k}}^{k})^{T}}||f_{k}\alpha^{k}||_{2}\quad\text{such that}\quad\sum_{i=0}^{m_{k}}\alpha_{i}^{k}=1;$
     Update $\mathbf{u}^{k+1}$ using Anderson mixing:  $\mathbf{u}^{k+1}=\sum_{i=0}^{m_{k}}\alpha_{i}^{k}P(\mathbf{u}^{k-m_{k}+i})$;
end for
5.1 Complexities/Costs of Fine-scale Methods
We now provide an overview of the complexity of each method. Recall that all proposed methods besides Newton’s method involve an NRAS update. Computing the NRAS update (16) requires solving a family of local nonlinear systems associated with the subdomains. The local nonlinear systems are computed by Newton’s method, with a linear system solved at each iteration. We denote by $N_{it,loc}$ the required number of local Newton iterations averaged over all subdomains. Due to the small size of the subdomains, the linearized system at each Newton iteration is computed via a direct sparse linear solve. This number might vary according to the method used, and we use an upper index to distinguish different case scenarios. We denote by $C(LU_{loc})$ a typical cost of the local LU decompositions on each subdomain, assumed to be approximately the same for every subdomain. Additionally, we denote by $C(\text{local assembly})$ the cost of assembling the local residual and Jacobian vectors, which will be done for all subdomains at each local iteration. The complexity of the NRAS iteration at each outer iteration can be expressed by:
$C(\text{NRAS})$ $=N\cdot N_{it,loc}\left(C(LU_{loc})+C(\text{local assembly})\right),$ (27)
where $N$ is the number of subdomains. Here, (27) assumes that the local computations are not done in parallel. With a parallel implementation, instead of the NRAS cost being summed over all subdomains, the cost would be dominated by the subdomain associated to the maximum number of local linear solves.
With the exception of Anderson’s method, all methods involve a global linear system which is solved via GMRES at each iteration. The complexity of the GMRES method for the Two-step and Newton methods at each outer iteration is given by the following:
$N\cdot C(LU_{loc})+C(LU_{coarse})+C(\text{assembly})+N_{\text{GM}}\cdot C(\text{GMRES}),$ (28)
where $N\cdot C(LU_{loc})+C(LU_{coarse})$ denotes the cost of assembling the preconditioner at each outer iteration, $C(\text{assembly})$ denotes the cost of assembling the global residual and Jacobian vectors/matrices $F({\bf u}),\nabla F({\bf u})$, and $C(\text{GMRES})$ denotes the cost of each GMRES iteration, which generally involves a sparse matrix-vector multiplication. We denote by $N_{\text{GM}}$ the average number of GMRES iterations per outer iteration. That is, $N_{\text{GM}}$ denotes the average number of GMRES iterations until convergence of the preconditioned linear systems occurring at each outer iteration, which will depend on the chosen solution strategy. We use the upper index to $N_{\text{GM}}^{i}$ to illustrate that this number can be different and decrease with the addition of an accelerator like a coarse space.
We note that in the RASPEN variants, the LU decompositions and assemblies for the nonlinearly preconditioned system can be reused from the earlier NRAS iteration, reducing the cost of the GMRES method compared to (28). That is, the complexity of the GMRES method for the One and Two-level RASPEN methods at each outer iteration is given by the following:
$N_{\text{GM}}\cdot C(\text{GMRES}),$
where this cost is equal to (28) for the Newton and Two-step methods. We denote by $C(\text{coarse})$ the cost of the sparse linear solve used for the coarse nonlinear problem in the Two-level RASPEN algorithm. Therefore, $N_{\text{c}}\cdot C(\text{coarse})$ denotes the total cost of the coarse nonlinear problems per outer iteration, where $N_{\text{c}}$ denotes the average number of coarse linear solves per outer iteration.
With this, the costs for each method are summarized to be
  • 1.
    Newton: $N\cdot C(LU_{loc})+C(LU_{coarse})+C(\text{assembly})+N_{\text{GM}}\cdot C(\text{GMRES})$;
  • 2.
    Two-Step: $C(\text{NRAS})+N\cdot C(LU_{loc})+C(LU_{coarse})+C(\text{assembly})+N_{\text{GM}}\cdot C(\text{GMRES})$;
  • 3.
    One-level RASPEN: $C(\text{NRAS})+N_{\text{GM}}\cdot C(\text{GMRES})$;
  • 4.
    Two-level RASPEN: $C(\text{NRAS})+N_{\text{GM}}\cdot C(\text{GMRES})+N_{c}\cdot C(\text{coarse})$;
  • 5.
    Anderson: $C(\text{NRAS})+C(\text{assembly})+C(\text{coarse})$.
Furthermore, we quantify the following costs as follows:
  • 1.
    $C(LU_{loc})=O(N_{loc}^{3/2})$, where $N_{loc}\approx\frac{N_{\Omega}}{N}$ denotes the number of degrees of freedom in a subdomain and is assumed to be roughly consistent over all subdomains;
  • 2.
    $C(LU_{coarse})=O(N_{\cal V}^{3/2})$;
  • 3.
    $C(\text{local assembly})=O(N_{loc})$;
  • 4.
    $C(\text{assembly})=O(N_{\Omega})$;
  • 5.
    $C(\text{GMRES})=O(N_{\Omega})$ for each sparse matrix-vector product;
  • 6.
    $C(\text{coarse})=O(N_{\cal V}^{3/2})$;
where the LU decomposition of some sparse $m\times m$ matrix is supposed to be $O(m^{3/2})$ [18].
We denote the total number of outer iterations by $N_{k}$; this number may vary according to the method used, and we can use an upper index to distinguish between methods. Finally, the total cost over all outer iterations is given by:
$C(\text{Newton})$ $=N_{k}^{0}\Bigl{(}$ $N$$\times$ $O(N_{loc}^{3/2})$ $+$ $(N_{\text{GM}}^{0}+1)$$\times$ $O(N_{\Omega})$ $+$ $O(N_{\cal V}^{3/2})\Bigr{)}$;
[…]
$C(\text{Anderson})$ $=N_{k}^{4}\Bigl{(}$ $N\times N_{it,loc}^{4}$$\times$ $O(N_{loc}^{3/2})$ $+$ $O(N_{\Omega})$ $+$ $O(N_{\cal V}^{3/2})\Bigr{)}$;
where the $O(N_{\Omega})$ term arises as the dominant term in the least-squares problem of Anderson’s method and $C(\text{assembly})$ is included in Anderson’s method as well. Additionally, we have simplified the cost of the local assembly as $N\cdot N_{it,loc}\cdot O(N_{loc})\approx N_{it,loc}\,O(N_{\Omega})$ from the definition of $N_{loc}$.
We remark that for a time-dependent problem such as (1), there will be an additional $N_{T}$ term multiplied into the cost of each method, where $N_{T}$ denotes the number of time steps. In this case, $N_{k}^{i}$ would denote the average number of outer iterations per time step. As shown, we can estimate the complexity of individual components of the methods such as the cost of assembly, the cost of a sparse linear solve, and the cost of a matrix-vector product. Additionally, we report the iteration counts such as $N_{k}^{i}$, $N_{\text{GM}}^{i},$ $N_{it,loc}^{i}$, and $N_{c}$ in Section 6.
For the Two-step, Newton, and Two-level RASPEN methods, the GMRES method occurring at each outer iteration will be preconditioned with some two-level DD preconditioner. Therefore, we anticipate that $N_{GM}^{0},N_{GM}^{1}$, and $N_{GM}^{3}$ will be approximately equal, with $N_{GM}^{2}$ for the One-level RASPEN method growing as we increase $N$. We also anticipate that $N_{it,loc}^{i}$ will be consistent between all methods which involve an NRAS solve (each method besides Newton). Therefore, the cost of the Two-step and Two-level RASPEN methods can be directly compared by comparing the the outer iterations $N_{k}$, where $N_{\cal V}<<N_{\Omega}$ such that the cost of a coarse linear solve will be significantly lower than a fine linear solve. For reference, we will still record the number of coarse linear iterations $N_{c}$ for the Two-level RASPEN to ensure it is not dominating the cost.
For Newton’s method, it is clear that the cost of each iteration is lower than the Two-step and RASPEN methods, as there is no NRAS solve involved. However, the main goal of the proposed methods is to substantially reduce the number of outer iterations to make up for the increased complexity within each outer iteration. That is, we expect $N_{k}$ to be reduced for the other methods, allowing for a potential gain in performance. With regard to Anderson’s method, it appears to have quite a low cost at each iteration, with one NRAS solve at each iteration and the cost of the least-squares problem dominated by one $O(N_{\Omega})$ term. However, we do expect a higher value of outer iterations for this method $N_{k}^{4}$ when compared to the Two-step and RASPEN variants. We note that while it is difficult to directly compare an outer iteration of Anderson to the other methods, it is still an interesting alternative that is worth exploring in this article.
With this, our main concern in terms of complexity lies in the term $N_{k}^{i}N_{\text{GM}}^{i}O(N_{\Omega})$, representing the total number of “inner” GMRES iterations. However, when applicable, we also wish to record the term $N_{k}^{i}N_{it,loc}^{i}$, representing the total number of sparse local linear solves on a given subdomain. In Section 6, we record the number of outer iterations $N_{k}^{i}$, the average number of GMRES iterations per outer iteration $N_{\text{GM}}^{i},$ the cumulative number of GMRES iterations $N_{k}^{i}N_{\text{GM}}^{i}$, as well as the average and cumulative (over all outer iterations) local linear solves per subdomain, namely $N_{it,loc}^{i}$ and $N_{k}^{i}N_{it,loc}^{i}$.
6 Numerical Results
We now provide numerical experiments for the proposed algorithms (Algorithms 1-5) for multiple test cases. These numerical tests will proceed in order of increasing difficulty; they include a stationary porous medium equation on an L-shaped domain, the similar equation on a small realistic urban domain, and the Diffusive Wave model on a large realistic urban domain.
For all numerical experiments, the overlap is set to $\frac{1}{20}\mathcal{H}_{j}$, where $\mathcal{H}_{j}=\max(x_{max,j}-x_{min,j},y_{max,j}-y_{min,j})$ and $x_{min,j}$, $y_{min,j}$, $x_{max,j}$, $y_{max,j}$ denote the minimal and maximal $x$ and $y$ coordinates that are contained in $\Omega_{j}$. We do not explore the effect of overlap further.
6.1 Porous Medium Equation on L-shaped Domain
As a first numerical example, referred to as Example 1, we use an L-shaped domain with a reentering corner. The domain is defined by $D=(-1,1)^{2}$, $\Omega_{S}=(0,1)^{2}$ and $\Omega=D\setminus\overline{\Omega_{S}}$ such that the model domain has a singularity/corner in the upper right quadrant of the domain. With this model domain, we consider a porous medium equation given by
$\left\{\begin{array}[]{rll}u(x,y)+\text{div}(\nabla(u(x,y)^{4}))&=&0\qquad\mbox{in}\qquad\Omega,\\ \displaystyle u(x,y)&=&1\qquad\mbox{on}\qquad y=1,\\ \displaystyle u(x,y)&=&0\qquad\mbox{on}\qquad x=1,\\ \displaystyle\frac{\partial u(x,y)}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad\Gamma_{N},\\ \end{array}\right.$ (29)
where $\Gamma_{N}=\{(x,y)\in\partial\Omega\,|\,x\neq 1\text{ and }y\neq 1\}$ denotes the Neumann boundary. Consider a piecewise affine finite element discretization, and let $\mathbf{u}$ denote a vector of interior node values of size $N_{\Omega}$. We denote the nodal “hat” basis functions $(\eta^{\ell})_{\ell=1}^{N_{\Omega}}$. We define the lumped finite-element mass matrix $M$ and finite element stiffness matrix $A$ such that
$M_{ii}=\sum_{\ell=1}^{N_{\Omega}}\int_{\Omega}\eta_{i}\eta_{\ell}\,\,{\rm d}\mathbf{x},\qquad\text{and}\qquad A_{i\ell}=\int_{\Omega}\nabla\eta_{i}\cdot\nabla\eta_{\ell}\,\,{\rm d}\mathbf{x}.$
With this, the matrix-vector notation of (29) becomes
$F(\mathbf{u})=M\mathbf{u}+A\max(\mathbf{u},0)^{4}.$ (30)
The finite element solution of this equation is shown in Figure 2.
Figure 2: Finite element solution of (29) for Example 1 on the chosen L-shaped domain.
Figure 3: Example 1, convergence curves. Top: $N=3$ (left), $N=5$ (right). Bottom: $N=9$.
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 88 38.1 3353 N/A N/A
Anderson 31 N/A N/A 4.35 134.85
2-step 22 35.0 770 3.96 87.12
1-level RASPEN 15 85.3 1280 3.7 55.5
2-level RASPEN 6 19.2(+5.3) 115(+32) 3.94 23.64
Table 1: Example 1, $N=9$, initial guess $\mathbf{u}^{0}=0.05$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 81 33.4 2705 N/A N/A
Anderson 32 N/A N/A 4.64 149.0
2-step 16 34.0 544 4.86 77.8
1-level RASPEN 11 53.9 593 5.08 55.9
2-level RASPEN 6 19.2(+4.5) 115(+27) 4.48 26.9
Table 2: Example 1, $N=5$, initial guess $\mathbf{u}^{0}=0.05$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 79 33.7 2662 N/A N/A
Anderson 25 N/A N/A 4.36 109.0
2-step 13 38.9 506 6.06 78.8
1-level RASPEN 9 35.4 319 6.96 62.6
2-level RASPEN 6 20.5(+3.8) 123(+23) 5.5 33.0
Table 3: Example 1, $N=3$, initial guess $\mathbf{u}^{0}=0.05$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
We provide numerical results for this example for Newton’s method given by Algorithm 1, the Two-step method from Algorithm 4, the One-level RASPEN method from Algorithm 2, the Two-level RASPEN method from Algorithm 3, and the Anderson acceleration method from Algorithm 5. For the Two-step and Newton methods, at each outer iteration, the linear system is solved using GMRES method combined with the linear two-level RAS preconditioner (14). We show results for coarse partitionings $N=9\times 9,5\times 5$, and $3\times 3$; The choice of $N$ being a square of an odd number ensures the consistency of the mesh sequence in terms of the shape of the elements.
Convergence curves for this example are shown in Figure 3. Additionally, results for this example are summarized in Tables 1, 2, and 3. These tables report the total number of outer iterations $N_{k}$ and, whenever appropriate, also report the average number of GMRES iterations $N_{\text{GM}}$, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$, the average number of local linear solves on a given subdomain per outer iteration $N_{it,loc}$, and the cumulative number of local linear solves on a given subdomain $N_{k}\cdot N_{it,loc}$.
In terms of outer iterations, from Figure 3, we see that Newton’s method results in a very large plateau before the region of quadratic convergence is reached. This difference is stark when compared to the other methods. Additionally, for all number of subdomains $N$, Two-level RASPEN outperforms the other methods in terms of outer iteration count, with the One-level RASPEN, Two-step, Anderson, and Newton methods following in order. Additionally, we notice that while Anderson does not initially plateau like Newton’s method, we do not obtain the eventual steep slope of convergence which the other methods achieve.
We see from Tables 1, 2, and 3 that the Two-level RASPEN method is scalable in terms of both outer iterations and GMRES iterations, in the sense that these values remain approximately the same for all $N$. In terms of GMRES iterations, we observe that the One-level RASPEN method loses scalability as the number of subdomains increases, more so than the other methods which all involve a coarse component of some type; particularly, as $N$ is doubled in one direction, the GMRES iterations increase and $N_{k}\cdot N_{\text{GM}}$ about doubles as well. We also observe that Newton’s method and the Two-step method, which both solve a linearized system with the linear two-level RAS preconditioner (with the Trefftz space as the coarse space), result in similar values of $N_{\text{GM}}$ for all $N$. However, as the number of Two-step outer iterations is lower than the number of Newton outer iterations, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$ is significantly lower for the Two-step method. The Two-level RASPEN method results in both the smallest value of $N_{k}$ and the smallest value of $N_{\text{GM}}$; therefore, it clearly results in the lowest number of cumulative GMRES iterations $N_{k}\cdot N_{\text{GM}}$. While the Two-level RASPEN has the additional $N_{k}\cdot N_{c}$ coarse linear solves, these coarse solves are generally less expensive than the global GMRES iterations, and the number of coarse linear solves is not exceedingly large.
In terms of local linear solves per subdomain, we do see a slight dependence of $N_{it,loc}$ on the number of subdomains $N$ for all methods, likely due to the fact that the size of the local problems $\frac{N_{\Omega}}{N}$ gets larger as $N$ is decreased. However, between methods for a given $N$, the values of $N_{it,loc}$ appear consistent as expected. Depending on the method, as $N_{k}$ is increased, clearly $N_{k}\cdot N_{it,loc}$ will increase accordingly.
6.2 Porous Medium Equation on Large Urban Domain
As a second example, referred to as Example 2, we provide numerical results for a Porous Medium equation on a large urban domain of size $160\times 160$ meters such that $D=(-80,80)^{2}$ and $\Omega_{S}$ is given by the union of realistic perforations. The buildings are removed from the computational domain, leading to 72 total perforations. The geometry of the buildings have been provided by Métropole Nice Côte d’Azur. After discretization, this model domain contains 35220 triangles and 21317 mesh points. As mentioned, the overlap is set to $\frac{1}{20}\mathcal{H}_{j}$.
For this example, we ensure that the fine-scale triangulation is identical as the number of subdomains $N$ changes by generating a finer background mesh to be used for various $N$; specifically, we can a fine-scale triangulation conforming to the coarse partitioning $N=16\times 16$ subdomains, then use this background triangulation for $N=2\times 2,4\times 4,$ and $8\times 8$ subdomains.
With this, we consider a porous medium equation given by
$\left\{\begin{array}[]{rll}u(x,y)+c\,\text{div}(\nabla(u(x,y)^{m}))&=&0\qquad\mbox{in}\qquad\Omega,\\ \displaystyle u(x,y)&=&c\qquad\mbox{on}\qquad x=-80,\\ \displaystyle u(x,y)&=&0\qquad\mbox{on}\qquad x=80,\\ \displaystyle\frac{\partial u(x,y)}{\partial\bf{n}}&=&0\qquad\mbox{on}\qquad\Gamma_{N},\\ \end{array}\right.$ (31)
where $\Gamma_{N}=\{(x,y)\in\partial\Omega\,|\,x\neq-80\text{ and }y\neq 80\}$ denotes the Neumann boundary; with this, the boundary of each perforation ($\partial\Omega\cap\partial\Omega_{S}$) is included in $\Gamma_{N}$. Here, we take $m=3$ and $c=15$ in (31). With a similar finite element discretization to Example 1, the matrix-vector notation of (31) becomes
$F({\bf u})=M{\bf u}+c\,A\max({\bf u},0)^{m},$
with the finite element solution for this example shown in Figure 4.
Figure 4: Finite element solution of (31) for Example 2 on the chosen urban model domain.
For this example, we provide numerical results for Newton’s method from Algorithm 1, the Two-step method from Algorithm 4, the Two-level RASPEN method from Algorithm 3, and the Anderson acceleration method from Algorithm 5. As we have established that One-level RASPEN is not a robust method for large $N$, we do not provide results for One-level RASPEN in this example. As in the previous example, for the Two-step and Newton methods, at each outer iteration, the linear system is solved using GMRES method combined with the linear two-level RAS preconditioner (14).
The results for this example are summarized in Tables 4, 5, 6, and 7 for initial guesses of $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$ and for subdomain partitioning of $N=16\times 16,8\times 8,4\times 4,$ and $2\times 2$, respectively. While the coarse partitionings vary, the background fine-scale triangulation is consistent throughout the numerical experiment. The total number of coarse nodes are given by $N_{\cal V}=$784, 332, 149, and 56 for $N=16\times 16,8\times 8,4\times 4,$ and $2\times 2$, respectively. These tables report the total number of outer iterations $N_{k}$ and, whenever appropriate, also report the average number of GMRES iterations $N_{\text{GM}}$, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$, the average number of local linear solves on a given subdomain per outer iteration $N_{it,loc}$, and the cumulative number of local linear solves on a given subdomain $N_{k}\cdot N_{it,loc}$.
Corresponding convergence curves for both initial guesses are provided in Figure 5. In terms of outer iterations, even more than the previous example, Newton’s method has a large plateau in convergence before reaching the region of quadratic convergence. For all number of subdomains $N$, Two-level RASPEN outperforms the other methods in terms of outer iteration count, with the Two-step, Anderson, and Newton’s methods following in increasing order. We particularly see that the outer iterations of the Two-level RASPEN method are extremely robust with respect to $N$, more so than the other accelerated methods. Additionally, from Tables 4, 5, 6, and 7, we see that there is a slight increase in $N_{k}$ for the Two-step and Anderson methods as $N$ is increased, but the increase in $N_{k}$ is not proportional to the increase in $N$. The number of outer iterations $N_{k}$ stays extremely consistent for the Newton and Two-level RASPEN methods as $N$ is varied. In terms of initial guess, we observe that Newton’s method is much more sensitive to the initial guess than the other methods in terms of outer iteration count. In fact, we see that this change in initial guess can result in a 50% increase in iteration count for Newton’s method. Other than the outer iterations for Newton’s method, there is no significant dependence on the initial guess.
From Tables 4, 5, 6, and 7, we see that $N_{\text{GM}}$ for the Two-level RASPEN method is significantly smaller than that of the Two-step and Newton methods. Additionally, as the number of outer iterations $N_{k}$ is the lowest for the Two-level RASPEN method, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$ is the lowest for the Two-level RASPEN method. As in the previous example, the Two-step method and Newton’s method result in around the same number of GMRES iterations per outer iteration $N_{\text{GM}}$, with this value slightly increasing as $N$ is increased; however, this increase is not proportional to the increase in $N$. As the Two-step method results in fewer outer iterations $N_{k}$ than Newton’s method, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$ is significantly lower for the Two-step method. This can result in between 10 and 20 times more total GMRES iterations for Newton’s method.
In terms of local linear solves per subdomain, our conclusions are similar to those of Example 1. Between methods, the values of $N_{it,loc}$ appear fairly consistent, but depending on $N_{k}$, $N_{k}\cdot N_{it,loc}$ will increase accordingly.
Overall, our conclusions from Example 2 are similar to that of Example 1, with Newton’s method particularly struggling in terms of outer iterations and having an additional dependency on the initial guess. However, in terms of solving the linearized system, it appears that the Trefftz coarse space is successful as a component of the linear two-level RAS preconditioner for this nonlinear problem.
Figure 5: Example 2, convergence curves. Solid lines correspond to an initial guess of $\mathbf{u}^{0}=1$, dotted lines correspond to an initial guess of $\mathbf{u}^{0}=0$. Top: $N=2\times 2$ (left), $N=4\times 4$ (right). Bottom: $N=8\times 8$ (left), $N=16\times 16$ (right).
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 45.0 10215 N/A N/A
Anderson 37 N/A N/A 4.1 152
2-step 25 41.1 1028 4.14 104
2-RASPEN 9 22.3(+7.4) 201(+67) 3.13 28
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 46.0 15916 N/A N/A
Anderson 52 N/A N/A 5.07 264
2-step 30 41.6 1248 2.83 85
2-RASPEN 8 19.2(+9.6) 154(+77) 2.11 17
Table 4: Example 2, $N=16\times 16$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 35.2 7990 N/A N/A
Anderson 35 N/A N/A 5.03 176
2-step 20 38.4 768 5.42 108
2-RASPEN 7 27.1(+8.0) 190(+56) 5.25 37
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 34.0 11764 N/A N/A
Anderson 35 N/A N/A 5.06 177
2-step 20 38.1 762 4.5 90
2-RASPEN 6 19.5(+8.5) 117(+51) 3.85 23
Table 5: Example 2, $N=8\times 8$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 26.2 5947 N/A N/A
Anderson 22 N/A N/A 7.66 169
2-step 11 30.1 331 9.19 101
2-RASPEN 7 20.4(+5.9) 143(+41) 7.87 55
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 25.4 8788 N/A N/A
Anderson 22 N/A N/A 7.96 175
2-step 11 31.9 351 9.66 106
2-RASPEN 7 15.3(+6.3) 107(+44) 6.07 43
Table 6: Example 2, $N=4\times 4$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 21.0 4767 N/A N/A
Anderson 17 N/A N/A 10.0 170
2-step 8 25.6 205 16.19 130
2-RASPEN 6 15.3(+4.3) 92(+26) 13.33 80
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 18.3 6332 N/A N/A
Anderson 17 N/A N/A 13.93 237
2-step 8 25.6 205 24.53 196
2-RASPEN 6 13.2(+4.3) 79(+26) 26.21 157
Table 7: Example 2, $N=2\times 2$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
6.3 Diffusive Wave Model on Large Realistic Urban Domain
In the final numerical experiment, referred to as Example 3, we consider the Diffusive Wave model (1) used to model a hypothetical flood in a densely urbanized area of the city of Nice, France. The model domain is depicted in Figure 8 and has dimension $850\times 1500$ meters. In this test case scenario, the flood is produced by an overflow of Paillon river in the north-west part of the domain. Once again, the buildings are removed from the computational domain, leading to numerous perforations (447 in total). The bathymetry $z_{b}$ uses the $1m$ Digital Elevation Model available from [20]. The parameters of the Diffusive Wave model are chosen formula as $\alpha=1.5$ and $\gamma=0.5$, based on Chézy’s formula, with friction coefficient $c_{f}=30$ which corresponding to a rough terrain [32]. We set Dirichlet boundary conditions, with water flow of about 2 meters coming from the Paillon river region and $u=z_{b}$ elsewhere; this boundary condition is shown in Figure 6. The initial condition is taken as equal to $z_{b}$ and the final simulation time is set to $T_{f}=1500$ seconds, which corresponds to a steady state solution.
Figure 6: Boundary condition (thick blue and red lines) and corresponding model domain for Example 3. Red corresponds to $u=21.5$ meters, blue corresponds to $u=z_{b}$. This results in around a 2-meter elevation for the portion of the boundary representing the Paillon river.
Similarly to Example 2, we ensure that the fine-scale triangulation is identical as the number of subdomains $N$ changes by generating a finer background mesh to be used for various $N$. We generate a fine-scale triangulation for $N=8\times 16$ subdomains, then use this background triangulation for $N=1\times 2,2\times 4,4\times 8$ and $8\times 16$ subdomain partitionings. Figure 7 shows the matching background triangulation for $N=2\times 4$ and $8\times 16$ subdomains, respectively.
Consider the NRAS iteration (16), which contains the solution of local nonlinear subproblems via Newton’s method. In practice, for time-dependent problems, we come across the issue that some local subproblems will not converge for a given time increment $\Delta t_{n}=t_{n+1}-t_{n}$ in the residual function (5). We have found for our numerical examples that often, it is a small portion of subproblems that will not converge, generally due to stagnation in convergence. Therefore, it may be inefficient to reduce the global time increment for all subproblems and the entire iteration. We implement a local time step reduction strategy which makes it possible to reduce the time increment locally, then use this solution as an initial guess for the original system. This process is described in Algorithm 3.
Procedure 3 Local subproblem with local time-stepping
$\Delta t,\mathbf{u}^{k}$, initial guess $G_{j}^{0}$, subdomain index $j$
Let $G_{j}(\mathbf{u}^{k};G_{j}^{0};\Delta t)$ denote the solution to the local subproblem (15) with initial guess $G_{j}^{0}$ and time step $\Delta t$ used in the residual.
$\Delta t_{worked}=\Delta t$
1: Solve local subproblem with some $\Delta t_{worked}\leq\Delta t$ for $G_{j}^{w}=G_{j}(\mathbf{u}^{k};G_{j}^{0};\Delta t_{worked}):$
return $G_{j}^{*}=G_{j}^{w}$ if $\Delta t_{worked}=\Delta t$
Initialize $\Delta t_{try}=\Delta t_{worked}$
if $\Delta t_{worked}<\Delta t$ then
     Until convergence:
     2: Use $G_{j}^{w}$ as initial guess for original $\Delta t$, try to solve for $G_{j}^{*}=G_{j}(\mathbf{u}^{k};G_{j}^{w};\Delta t)$
     if  Step 2 fails: Initial guess was insufficient to solve original system then
         Form better initial guess: Increase $\Delta t_{try}=\frac{\Delta t+\Delta t_{try}}{2}$ and proceed to step 3
     else Return $G_{j}^{*}$
     end if
     3: Obtain new initial guess: Reset and solve for $G_{j}^{w}=G_{j}(\mathbf{u}^{k};G_{j}^{w};\Delta t_{try})$
     if  Step 3 fails: Solve didn’t converge for $\Delta t_{try}$ then
         Reduce $\Delta t_{try}=\frac{\Delta t_{worked}+\Delta t_{try}}{2}$ and return to step 3
     else Solve converged for $\Delta t_{try}$ and new initial guess $G_{j}^{w}$ obtained
         Set $\Delta t_{worked}=\Delta t_{try}$ and return to step 2
end if
end if
In this numerical experiment, except for Newton’s method, the discussed local adaptive time-stepping method of Algorithm 3 is implemented with a global time increment of $\Delta t=10$ seconds for each time step for all methods. We find that with this local adaptive time stepping, a global time increment reduction is not necessary for convergence.
For Newton’s method, we begin with an initial time increment of $\Delta t_{0}=10$ seconds. If the system at a given time step can not be solved, the global time increment will be reduced by a factor of $\sqrt{2}$ until the reduced system is able to be solved. Then, once the reduced system is solved, we increase the time increment by a factor of $\sqrt{2}$ for the next time step, with a maximum time increment of 10 seconds. This means that $\Delta t_{n+1}=\min(\sqrt{2}\Delta t_{n},10)$. As mentioned, the final time for this experiment is $T_{f}=1500$.
Figure 8 reports the simulation results for time $t=10,250,500,$ and 1500 seconds. In particular, in color, we report the water depth $\mathbf{h}=\mathbf{u}-z_{b}(\mathbf{x})$ over the “flooded region” where $\mathbf{h}\geq 1$cm. The ground surface elevation $z_{b}$ is reported in black and white.
The spatial discretization results in 81444 mesh points and 131581 triangles in the model domain. As mentioned, the fine-scale triangulation will be conforming to the polygonal coarse partitioning. Although we record results for various numbers of subdomains $N$, the fine-scale mesh will be kept consistent throughout the experiment and is conforming to a $N=8\times 16$ coarse mesh; therefore, it will be conforming to $N=1\times 2$, $N=2\times 4$, $N=4\times 8$ meshes as well. The coarse and fine partitionings are shown in Figure 7 for multiple coarse partitionings. By the end of the simulation, 38216 points and 60846 triangles are “flooded” such that $h\geq 1$cm. The total number of coarse nodes are given by $N_{\cal V}=$795, 357, 126, and 34 for $N=8\times 16,4\times 8,2\times 4,$ and $1\times 2$ subdomains, respectively. The overlap is once again set to $\frac{1}{20}\mathcal{H}_{j}$.
Figure 7: Coarse (thick blue lines) and fine (thin blue lines) discretizations for $N=2\times 4$ (left) and $N=8\times 16$ (right) subdomains for model domain representing the Nice port and Paillon river.
Recall that at every time step, the nonlinear system (5) has to be solved. For this example, in addition to Newton’s method given by Algorithm 1, we provide numerical results for the One-level RASPEN method from Algorithm 2, the Two-level RASPEN method from Algorithm 3, the Two-step method from Algorithm 4, and the Anderson acceleration method from Algorithm 5.
Figure 9 displays the cumulative number of (successful) outer iterations over the simulation time, obtained for each of the methods for various numbers of subdomains $N$.
In addition, we record various values in Tables 8, 9, 10 and 11. For this time-dependent example, these values report the total number of time steps $N_{T}$, the total number of outer iterations over time $N_{T}\cdot N_{k}$, and, whenever appropriate, also report the cumulative number of GMRES iterations over time $N_{T}\cdot N_{k}\cdot N_{\text{GM}}$ and the total number of local linear solves on a given subdomain over all time steps and outer iterations $N_{T}\cdot N_{k}\cdot N_{it,loc}$. For the Two-step and Newton methods, at each outer iteration, the linear system is solved using GMRES method combined with the linear two-level RAS preconditioner (14).
Additionally, we report the number of GMRES iterations per outer iteration $N_{\text{GM}}$ for each time step in Figure 10. As mentioned, the overall (cumulative) number of GMRES iterations is reported in Tables 8, 9, 10, and 11 for $N=8\times 16,4\times 8,2\times 4,1\times 2$, respectively.
Figure 8: Numerical solution of Example 3 at various time steps, where the numerical solution is presented by $\mathbf{h}=\mathbf{u}-z_{b}(\mathbf{x})$. Top: $t=10$ (left) seconds, $t=250$ (right) seconds. Bottom: $t=500$ (left) seconds, final $t=1500$ (right) seconds. “Flooded” region (where $\mathbf{h}\geq 1$cm) are shown in color, with underlying bathymetry $z_{b}$ shown in the background in black and white.
Figure 9: Example 3, cumulative outer iterations over time. Top: $N=1\times 2$ (left), $N=2\times 4$ (right). Bottom: $N=4\times 8$ (left), $N=8\times 16$ (right).
Figure 10: Example 3, average number of GMRES iterations per outer iteration $N_{\text{GM}}$ at each time step. Top: $N=1\times 2$ (left), $N=2\times 4$ (right). Bottom: $N=4\times 8$ (left), $N=8\times 16$ (right).
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
Newton 155 1797 44146 N/A
Anderson 150 2524 N/A 6849
2-step 150 687 20029 1557
1-level RASPEN 150 686 100453 1574
2-level RASPEN 150 730 12117(+2907) 2017
Table 8: Example 3, $N=8\times 16$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
Newton 160 1807 33432 N/A
Anderson 150 1902 N/A 6967
2-step 150 638 13164 1827
1-level RASPEN 150 624 33117 1824
2-level RASPEN 150 650 8065(+2031) 2334
Table 9: Example 3, $N=4\times 8$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
Newton 151 1776 26753 N/A
Anderson 150 1657 N/A 8538
2-step 150 592 9483 2412
1-level RASPEN 150 571 9040 2478
2-level RASPEN 150 595 6328(+1755) 3066
Table 10: Example 3, $N=2\times 4$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
Method $N_{T}$ $N_{T}\cdot N_{k}$ $(N_{T}\cdot N_{k}\cdot N_{\text{GM}})$ $N_{T}\cdot N_{k}\cdot N_{it,loc}$
Newton 155 1791 21608 N/A
Anderson 150 1225 N/A 7132
2-step 150 597 7794 3831
1-level RASPEN 150 477 3827 3377
2-level RASPEN 150 515 3983(+1114) 3698
Table 11: Example 3, $N=1\times 2$ subdomains. The value in brackets represents $N_{t}\cdot N_{k}\cdot N_{c}$.
From Figure 9, we see that the RASPEN variants (both one and two-level) as well as the Two-step method produce the smallest number of total outer iterations, and are very similar to one another in this regard. Additionally, we see that the Anderson acceleration changes the most as we increase $N$, with the outer iterations generally increasing. As expected, Newton’s method results in a higher number of outer iterations when compared to the RASPEN and Two-step methods.
We see from Figure 10 that as expected, as the number of subdomains is increased, the One-level RASPEN method is no longer scalable in terms of GMRES iterations; in fact, as $N$ grows, the average GMRES iterations $N_{\text{GM}}$ are significantly larger than the other methods such that they are out of the scope of the vertical axis in the figure (green dots in Figure 10). Additionally, as the number of subdomains grow, the Two-level RASPEN method clearly becomes the method with the lowest value $N_{\text{GM}}$ per outer iteration over time. As in the previous examples, for Newton’s method preconditioned with the two-level linear RAS preconditioner (blue dots in Figure 10) and the Two-step method preconditioned with the two-level linear RAS preconditioner (orange dots in Figure 10), the number of GMRES iterations are not increasing proportionally with $N$. However, we do see a slight increase as the number of subdomains is increased. We additionally remark that over time, the average number of GMRES iterations per outer iteration $N_{\text{GM}}$ in Figure 10 increases for each method besides the Two-level RASPEN method; this is generally most likely due to an increased proportion of the model domain being “flooded”. That is, the problem becomes increasingly difficult over time. However, this also suggests that the performance of the coarse correction is improved when applied “nonlinearly”, as is done in the Two-level RASPEN method.
From Table 8, 9, 10, and 11, we see the results from Figures 9 and 10 summarized. Particularly, Newton’s method results in much larger numbers of outer iterations when compared to the nonlinearly preconditioned variants (with the exception of the Anderson acceleration as $N$ grows large). Additionally, the total number of GMRES iterations $N_{T}\cdot N_{k}\cdot N_{\text{GM}}$ in the One-level RASPEN method grows significantly as $N$ is increased, more than doubling as $N$ is doubled in one direction. It appears that particularly as the number of subdomains grows, Two-level RASPEN is the most efficient, producing the lowest number of total GMRES iterations with a relatively small number of coarse linear solves. We note again that the coarse linear solves are computationally less expensive than a global fine-scale linear solve. However, we do remark that we do not see the same type of acceleration in terms of outer iterations for Two-level RASPEN like we obtained in the other two numerical examples. Overall, particularly if one is choosing a high value of $N$, it appears that the Two-level RASPEN method outperforms the others.
We have mentioned that from Figure 9, the outer iterations of the Anderson method appear to have noticeably more dependence on $N$ than the outer iterations of the other methods. To view this phenomenon, we provide the average outer iterations per time step $N_{k}$ for Anderson’s method in Figure 11. We see from Figure 11 that the outer iterations do increase noticeably with $N$. But in fact, the number of outer iterations stays reasonably stable as time increases, and the pattern of outer iterations seems to match that of average GMRES iterations of two-level RASPEN in Figure 10. This dependence on $N$ suggests that the coarse space could be improved for this model problem.
Figure 11: Example 3, Anderson outer iterations per time step. Colors correspond to various coarse partitionings.
7 Summary and Future Work
This article has provided multiple efficient nonlinear domain decomposition methods to solve nonlinear equations on complex, perforated domains. Specifically, we have focused on a Diffusive Wave model on a domain with numerous perforations; these perforations represent buildings and walls/fences in urban areas. The coarse space form [4], designed for linear problems on perforated domains, proves to form efficient two-level methods for our difficult model problem.
From our numerical experiments, we draw the following conclusions. In terms of the number of GMRES iterations, as expected, the One-level RASPEN method is not a robust method as the number of subdomains $N$ grows large. For the linearized systems resulting at each iteration of the Two-step and Newton methods, the Trefftz space performs well as a component of the two-level RAS preconditioner and results in a reasonable number of inner GMRES iterations per outer iteration. However, these iterations slightly increase as $N$ is increased. This effect could be mitigated by updating the basis functions in the Trefftz space as the linearized system is updated. The Two-level RASPEN method further reduces the total number of GMRES iterations compared to the Two-step method, particularly as $N$ grows. However, in terms of total GMRES iterations $N_{k}\cdot N_{\text{GM}}$, due to the large number of $N_{k}$ for Newton’s method, Newton’s method results overall in a much larger total number of GMRES iterations.
In terms of total outer iterations, Newton’s method is much more sensitive to the initial guess than the other proposed methods in our experiment, with iteration counts varying by up to 50%. Additionally, Newton’s method generally results in extremely large iteration counts when compared to the Two-step and RASPEN variants. The addition of an NRAS solve in the Two-step method greatly accelerates convergence in terms of outer iterations when compared to Newton’s method. Specifically, for the stationary examples, the Two-level RASPEN method provides a significant acceleration in terms of outer iterations when compared to the other proposed methods. We do not see the same effect in the time-dependent example, likely due to the time stepping process and choice of the time increment. The Anderson acceleration of a coarse Two-step method is also an interesting alternative to Newton’s method, reducing the total iteration count, particularly for the stationary examples. However, for the time-dependent example, we do see a dependence on $N$ in terms of outer iterations for this method. Each iteration of Anderson’s method involves a small least-squares problem and is cheaper than a Newton iteration. However, it does not appear to accelerate convergence when compared to methods such as the Two-level RASPEN method.
In terms of implementation, both Newton’s method and the Two-step method involve a sparse linear system at each iteration which, in principle, could be solved via a direct solve (provided the system is not too large). Additionally, the Two-step method is quite straightforward to implement when compared to the Two-level RASPEN method, with the only addition from Newton’s method being a nonlinear RAS computation at each iteration. The Two-level RASPEN method involves a nonlinear RAS computation, a coarse nonlinear problem, and a global linear system which generally should be solved via a Krylov method. However, once implemented, the Two-level RASPEN method forms an extremely robust and efficient algorithm. Additionally, Anderson acceleration benefits from readily available built-in functions provided by multiple popular programming languages.
Finally, we have found that local time step reduction is an efficient method to eliminate the need for a global time step reduction for all subdomains. Future work involves a full parallel implementation to provide detailed CPU times.
Acknowledgments
This work has been supported by ANR Project Top-up (ANR-20-CE46-0005). The high-resolution structural data has been provided by Métropole Nice Côte d’Azur. We warmly thank Florient Largeron, chief of MNCA’s SIG 3D project, for his help in preparation of the data and for the multiple fruitful discussions.